Calculus II, Spring 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 03

1. Section 9.5: Problems 4, 8, 7, 9, 24, 25, 33, 35, 37, 43, 45, 47.

2. Section 9.6: Problems 11, 15, 21, 23, 27, 28, 35, 39, 41, 44, 47.

3. (s9.6-extra1)

Find a power series whose interval of convergence is [1, 3). Do the same for (1, 3), [1, 3] and (1, 3], respectively.